Given the formula A=πr2, where A=1256 m2 and π=3.14/ descrbe the steps used to solve for "r". Showing is not enough
Answer by nyc_function(2626)

Given the formula A = πr^2, where A=1256 m^2 and π =3.14/ descrbe the steps used to solve for "r".
1256 = (3.14)(r^2)
Steps:
(1) divide both sides by 3.14.
(2) take square root of both sides.

1256/3.14 = r^2
400 = r^2
sqrt{400} = sqrt{r^2}
20 = r